kZAvala:

Angles A and B are complementary. What is the value of x? 34 56 68 79

Shadow:

Picture

Shadow:

Need the image

kZAvala:

1 attachment
Shadow:

Do you know the definition of a complementary angle?

Shadow:

or rather, if two angles are complementary, what does that mean?

kZAvala:

when two angles equal to 90

Shadow:

So if A and B are complementary, what does that mean?

kZAvala:

that they both equal to 90

Shadow:

So in order to solve for x, what could you do?

kZAvala:

subtract 22 from 90

Shadow:

Complementary angles: the sum of two angles which is 90 Sum means the result of an addition operation. This means that: \[A + B = 90\]

Shadow:

What can you do then?

kZAvala:

subtract?

Shadow:

Subtract what?

kZAvala:

90 from 22

Shadow:

Well, we have the values for angle A and B, so we can input them into the equation I made. \[A + B = 90\] \[(x + 22) + x = 90\] \[x + 22 + x = 90\] \[2x + 22 = 90\]

Shadow:

Does that make sense to you?

kZAvala:

yees yees now it does

kZAvala:

thank you

Shadow:

What did you get for x?

Shadow:

Well, glad you get it now. I'm going to close this question since it's been answered. Let me know if you have any further questions.